# Deep learning inference of the neutron star equation of state

### v1

@@ III.2 Deterministic deep network
$l_{\tt bin\_crossentropy}=-\frac{1}{N}\sum_{i=1}^{N}[y_{i}\cdot\text{log}(p_{i})\\ +(1-y_{i})\cdot\text{log}(1-p_{i})],$ (15)
@@
where $y$ is the actual class label, and $p_{i}$ is the predicted probability of the data belonging to the class. We also choose binary_accuracy as metric. We trained each model for $100$ epochs, and we found that a batch size of $100$ is optimal for most configurations. However, for most of the cases where $n=100$, a batch size of $50$ was preferred. We summarize the most efficient models we trained in tables I and II in the supplementary material available at Ref. [48], for the $M$-$R$ input model and the model with $M$-$R$-$k_{2}$ respectively. Since we do not fix the random-number seed of the network, each time the same model is trained from the start it exhibits slightly different loss and accuracy at the end of training. Hence, we trained each network $20$ times, and report the mean binary accuracy and validation binary accuracy with their respective standard deviation. We can see that the model accuracy increases with the number of hidden layers, the number of nodes, the number of noise injections ($n$) and the size of the observation set ($N$). However, for high values of number of noise injections and large observations set, training networks with more than a single hidden layers and a large number of nodes is computationally demanding. For the $M$-$R$ data, the best accuracy we can reach is $0.889\pm 0.001$ with 2 hidden layers of 215 and 200 nodes respectively, for a dataset with $n=300$ and $N=30$; while for the $M$-$R$-$k_{2}$ the best accuracy we obtain is $0.878\pm 0.003$ for a single hidden layer network with 270 nodes, with the dataset corresponding to $n=300$ and $N=30$.
@@
Let us now focus on the regression model, which aims to determine the remaining 15 parameters of the EOS. As in the previous case, we tested different neural networks varying the same parameters as above. We developed our models using the Python library Keras with TensorFlow. For every network we tested, we only allowed for a single hidden layer, while varying its number of nodes. We have seen that including more layers does not improve the efficiency of the model, while it requires more computational time. We also chose the ReLU activation function, with the He normal distribution as kernel initializer. The output layer consists of 15 nodes, corresponding to the 15 EOS parameters (7 $\rho$, 7 $c_{s}$, and $\Lambda$), together with the sigmoid activation function. The latter ensures that the speed of sound is bounded in the causal range $c_{s}\in(0,1)$  [^3] Note that, as a result of the renormalization process discussed in Sec. III.1, all parameters, not only the speed of sound, are bounded in the interval $(0,1)$.. We use the Adam fitting algorithm [67], and we employed the msle loss function, that is
$l_{\tt msle}(y,\hat{y})=\frac{1}{N}\sum_{i=1}^{N}[\text{log}(1+y_{i})-\text{log}(1+\hat{y}_{i})]^{2},$ (16)
@@
Figure 2: A specific random sample for the speed-of-sound vs normalized mass-density parametrization, where $\rho_{f}=12\,\rho_{0}$ is the last fixed value for the mass density. The black line describes the real model, while blue curves represent all predicted profiles by the regression deep network applied on test dataset augmented by injecting Gaussian noise $n=100$ times, following the procedure described in Sec. III.1. We also show the $2\,\sigma$ confidence interval for the network predictions as a blue shaded region, obtained by computing the mean and the standard deviation of the predictions at each value of renormalized mass density. The network predictions prefers to ‘settle’ around a mean value of $c_{s}$ such that $\langle c_{s}\rangle\sim 0.5$ for $\rho/\rho_{f}\to 1$. This behavior is retrieved generically, and is not related to the specific speed-of-sound model we show here.
@@
Figure 3: Predicted mean speed-of-sound $\langle c_{s}\rangle$ with a $3\,\sigma$ confidence interval as a function of the real mean value for $c_{s}$ for all possible different speed-of-sound realizations in the test dataset. Almost all predicted values are not in agreement with the expected ones. The deep network tends to smooth out the sound speed profile, and misses large increases or decreases in the profile, which are responsible for oscillations on the mean value of $c_{s}$. Indeed, one can notice that the former are bound in the interval $(0.5,0.6)$, whereas for the latter one finds $\langle c_{s}\rangle_{\text{real}}\in(0.4,0.8)$. One can notice that deviations increase at the boundaries of such interval, where the mean value of the real speed-of-sound is either lower or higher to the predicted ones due to the presence of large oscillations in the real profile not reproduced by the network’s realization.
Let us now focus on the results for the predicted vacuum energy shift. In Fig. 4, we show the predicted mean $\Lambda$ with a $2\,\sigma$ confidence interval as a function of the real one for all $10$ possible values considered. We considered here the $n=100$ noise injections dataset. Almost all predicted values lie within the $2\,\sigma$ intervals, with the exception of very large positive values of $\Lambda$. In particular, we notice that the highest possible value for the vacuum energy shift, i.e. $\Lambda=(194\,\text{MeV})^{4}$, is not correctly predicted, even when considering a $3\,\sigma$ interval. We expect that the reason why our networks can better identify a large negative phase transition, rather than a (large) positive one, lies on the different properties of these two cases. EOSs with a vacuum energy phase transition can be divided into two classes [46]: EOSs that admits a wide range of values for $\Lambda$ and EOSs that are allowed only if a large negative jump in vacuum energy is introduced. These lead to different $M$-$R$ and $k_{2}$-$R$ relation properties. In particular, EOSs that admit a large negative $\Lambda$ display peculiar $M$-$R$ relations and the network can better distinguish them between all possible configurations. Moreover, our training dataset is biased towards lower values of the vacuum energy shift, since these values yield a larger set of physically acceptable solutions of the gravity equations.
@@
Figure 5: Two specific EOS models (black lines) and their predictions for the test dataset augmented by injecting Gaussian noise $n=10$ times, following the procedure described in Sec. III.1. On the left (right) panel we show an EOS built from the AP4 (SLy) and with a negative (positive) jump in vacuum energy, with $\Lambda=-(194\,\text{MeV})^{4}$ ($\Lambda=(120\,\text{MeV})^{4}$). The two cases also use different speed-of-sound parametrizations. In almost all cases, the realizations have been predicted with the correct low-density model. At higher density scales, for the case of a negative vacuum energy transition, all predictions correctly display a sudden decrease in energy density. For a positive jump in $\Lambda$, however, the typical phase transition behavior is lost, and the energy density only slowly increases. The reason why our network can better identify a large negative vacuum energy, rather than a (large) positive one, lies on the different properties of these two cases. As it was pointed out in Ref. [46], there are two families of EOSs: those that admits a wide range of values for $\Lambda$ and those that are allowed only if a large negative jump in vacuum energy is introduced. This leads to different stellar properties. In particular, the second family displays peculiar $M$-$R$ relations and the network can better distinguish them between all possible configurations. Note also that our training dataset is biased towards lower values of the vacuum energy jump, since these choices lead to a larger set of physically acceptable solutions of the gravity equations.
@@
We now focus on the results coming from both the classification and the regression networks. This allows us to construct the full EOS, composed by the three regions described in section II. In Fig. 5, we show two specific EOS models (black line) and their predictions for the $n=10$ test dataset. Note that the two expected EOSs correspond to different low density descriptions, different speed-of-sound parametrization and different vacuum energy shifts. On the left panel we present an EOS built from the AP4 and with a negative jump in vacuum energy, with $\Lambda=-(194\,\text{MeV})^{4}$. For the low density region, $8$ out of $10$ realizations have been predicted with the correct model (AP4). Moreover, at higher density scales all predictions correctly follow the decrease in energy density due to the vacuum energy shifts. On the right panel we show the case of an EOS built from the SLy model with a positive vacuum energy jump, with $\Lambda=(120\,\text{MeV})^{4}$. In this instance, only a single realization is predicted with the wrong low-density description. However, at high densities, the predictions follow the expected behavior less rigorously, in particular the increase in energy density caused by the positive jump in vacuum energy is less abrupt, and one loses the typical phase transition behavior. Once again, the reason why our networks can better identify a large negative phase transition, rather than a (large) positive one, lies on our training dataset being biased towards lower values of the vacuum energy shift and on the fact that these two cases belong to different EOS’ families, which leads to different stellar properties [46].
@@
This can be seen in Fig. 6, where we show the $M$-$R$ relations obtained by solving the TOV equations for both the expected EOS (black line) and the predicted ones for the two EOS models considered in Fig. 5. Note that the color mapping is maintained, i.e. an EOS described by a specific color line in Fig. 5 yields a $M$-$R$ curve of the same color in Fig. 6. We also include, as an example, a single set of observation points $O=\{M_{i},R_{i}\}$ with injected Gaussian noise as described in Eq. (14) with a $1\,\sigma$ confidence interval. For both cases presented, all of the observation points, which we remind are used as input for our networks, are within $3\,\sigma$ from the $M$-$R$ curves obtained from the predicted realizations. This is true even when the model predicts a wrong low-density model.
@@
Figure 6: Mass-radius relations for two specific EOS models, corresponding to those presented in Fig. 5. On the left (right) panel we show an EOS built from the AP4 (SLy) and with a negative (positive) jump in vacuum energy, with $\Lambda=-(194\,\text{MeV})^{4}$ ($\Lambda=(120\,\text{MeV})^{4}$). Curves obtained from solving the TOV equations for the real EOS are shown in black, whereas the network predictions are shown with different colors, maintaining the same mapping as in Fig. 5. We also include, as an example, a single set of observation points with injected Gaussian noise as described with a $1\,\sigma$ confidence interval, as described in Sec. III.1, used as an input for the network. All of the observation points are within $3\,\sigma$ from the $M$-$R$ curves obtained from the predicted realizations.
@@
Finally, we discuss the $k_{2}$-$M$ relations obtained by solving Eq. 7 for the expected EOS (black line) and those obtained from our network. Once again, we maintain the color mapping as in Fig. 5. We do not include a set of observation points, to not complicate further the figure, however all mock observations are in agreement with the results produced from the predicted realizations. In Fig. 8, we show the fractional error for the tidal Love number $(k_{2}-k_{2,\,\text{real}})/k_{2,\,\text{real}}$. We can see that, typically, this is around $10\%$ at intermediate masses, while it can increase up to $30\%$ in the case of more massive stars.
@@
Figure 8: Fractional error for the tidal Love number as a function of the total stellar mass for two specific EOS models, corresponding to those presented in Fig. 5. On the top (bottom) panel we show an EOS built from the AP4 (SLy) and with a negative (positive) jump in vacuum energy, with $\Lambda=-(194\,\text{MeV})^{4}$ ($\Lambda=(120\,\text{MeV})^{4}$). We maintain the color mapping as in Fig. 5. The fractional error is typically around $10\%$ at intermediate masses, while it can increase up to $30\%$ in the case of more massive stars. This is to be expected since more massive configurations have higher central pressure, where the EOSs predictions deviates more significantly from the expected one, as can be seen from Fig. 5.
III.3 Bayesian deep network
Figure 9: A statistical sample of the predictions (features) as outputted from our Bayesian (probabilistic) deep network, and for two input choices according to our dataset’s numbering, corresponding to the $100$-th (top) and $1000$-th (bottom) input data point respectively. For a given input of masses, radii and tidal Love numbers, or equivalently a given input of the EOS, the output of the deep network corresponds to a particular realisation of the $7$ pairs of mass density-sound speed. We use the dataset where each input data point comprises of $30$ triplets of total mass, radius, tidal number, and we consider an $n_{\text{samples}}=100$ realisations of the output (coloured points). The red crosses denote the expected (“real”) values on the mass density-sound speed plane for the given EOS. The black dots and the associated error-bars are computed as the average and $2\sigma$ errors respectively, given the statistical sample for each column of the output’s array. As long as the predictions (red crosses) lie within the respective $2\sigma$ error bar, the prediction is considered successful, according to our evaluation pipeline, explained in Sec. III.3. Notice that in the plots we chose not to discard predictions which are outside the physical range for the mass and sound speed we consider, i.e points lying outside the range $[0,1]$. Though individual predictions may lie outside this range, central values and respective error bars are consistent with our assumed physical range of values. It is also to be noticed how the size of error bars change with the assumed EOS provided as input to the deep network.
@@ III.3 Bayesian deep network
$p({\boldsymbol{\theta}}|D)=\frac{p(\boldsymbol{\theta})\cdot p(D|\boldsymbol{\theta})}{\sum_{\boldsymbol{\theta}}p(D|\boldsymbol{\theta})\cdot p(\boldsymbol{\theta})}\equiv\frac{p(\boldsymbol{\theta},D)}{p(D)},$ (17)
@@
$q({\boldsymbol{\theta}})\approx p({\boldsymbol{\theta}}|D).$ (18)
@@
$\text{KL}\left(q||p\right)\equiv\sum_{\boldsymbol{\theta}}q(\boldsymbol{\theta})\log\frac{q(\boldsymbol{\theta})}{p(\boldsymbol{\theta}|D)}.$ (19)
In our approach, the approximate posterior, $q(\boldsymbol{\theta})$, will be based on some analytical, a priori ansatz, which we will compute after minimisation of the KL divergence (19). The resulting posterior function will then be a proxy to the true one in Bayes’ rule  [^4] We notice here that the KL divergence extends to the case of a continuous random variable through the replacement of the sum with an integral.. It can be shown that minimising the KL divergence amounts to maximising the quantity
$\mathds{E}_{q}\left[\log p(\boldsymbol{\theta},D)-\log q(\boldsymbol{\theta})\right]=\log p(\boldsymbol{\theta})-\text{KL}\left(q||p\right),$ (20)
@@
We implement our Bayesian deep network on TensorFlow using the class of DenseVariational layers which is part of the tensorflow_probability library. These are the probabilistic extension of the more standard Dense layers. We choose a LeakyReLU for our activation function and the mean-square error for our loss function respectively. The activation function is applied to all hidden layers, but not in the output layer - we comment further below about this choice. The minimisation of the loss function is performed through the RMSprop algorithm, which shows good-enough convergence at training. We found that a network with $4$ hidden layers, each with $16$ units, was a sufficiently minimal choice yielding satisfactory results for our purposes. Notice that the input and output of the network remained the same as with the standard (non-probabilistic) version of the deep network.
@@
$\bar{\rho}^{(J)}=\frac{1}{N}\sum_{i=1}^{i=N}\rho_{i}^{(J)},\;\;\;\;\bar{c}_{s}^{(J)}=\frac{1}{N}\sum_{i=1}^{i=N}c_{s\,i}^{(J)},\;\;\;J=1,\ldots,7,$ (21)
@@
$\bar{c_{s}}^{(J)}-2\sigma_{c_{s}}^{(J)}\leq c_{s\,\text{true}}^{(J)}\leq\bar{c_{s}}^{(J)}+2\sigma_{c_{s}}^{(J)}$ (22)
An example result of this process for a given data input is shown in Fig. 9 and 11 for a specific choice of input. For each data point we define as the accuracy $(\text{acc}_{i})\equiv N_{\text{accepted}}/7$, and the total accuracy as $\text{acc}\equiv\sum_{i}^{N}(\text{acc})_{i}/N,$ with $N$ the number of data points. The choice of the $2\sigma$ range for accepting/rejecting points is clearly a choice that we make as a reasonable credibility interval.
@@
$\text{mean accuracy}\pm\sigma\simeq\left(79\pm 16\right)\%.$ (23)
The spread of $16\%$ around the mean accuracy is somewhat large, implying that our model performs better on some EOS than others. It is to be emphasised that the features of our training/testing dataset, i.e the $7$ values of the sound speed, were drawn randomly from a normal distribution, as explained in Sec. III.1. Our probabilistic network, therefore, did not have sufficient information to reconstruct the posterior functions more accurately and in turn, achieve a better accuracy. We also note that the accuracy of the prediction decreases with decreasing the number of input observations in the network, i.e the number of stellar mass, stellar radius and tidal Love number triplets. This we have verified trying with a dataset where each input to the network consisted of $15$ and $30$ pairs respectively. It is to be further noted that, the accuracy of the network’s predictions on test data may increase with decreasing training accuracy. This is because the error bars of output statistical samples will be larger at lower training accuracy, a situation which is clearly undesirable. We found that an optimal choice for the training accuracy is around $70\%$, which is a fair compromise between the previous issue, and the also undesirable problem of overfitting. At this point, we should emphasize that the quoted training accuracy is according to the way that TensorFlow translates the value of the loss function to an accuracy value. In our Bayesian setting, however, the accuracy cannot be derived by simply comparing the network’s single prediction with the true values of the features, but through a statistical analysis of the respective output samples.

### v2

@@ 3.2 Deterministic deep network
$l_{\tt bin\_crossentropy}=-\frac{1}{N}\sum_{i=1}^{N}[y_{i}\cdot\text{log}(p_{i})+(1-y_{i})\cdot\text{log}(1-p_{i})],$ (3.3)
@@
where $y$ is the actual class label, and $p_{i}$ is the predicted probability of the data belonging to the class. We also choose binary_accuracy as metric. We trained each model for $100$ epochs, and we found that a batch size of $100$ is optimal for most configurations. However, for most of the cases where $n=100$, a batch size of $50$ was preferred. We summarize the most efficient models we trained in tables I and II in the supplementary material available at Ref. [51], for the $M$-$R$ input model and the model with $M$-$R$-$k_{2}$ respectively. Since we do not fix the random-number seed of the network, each time the same model is trained from the start it exhibits slightly different loss and accuracy at the end of training. Hence, we trained each network $20$ times, and report the mean binary accuracy and validation binary accuracy with their respective standard deviation. We can see that the model accuracy increases with the number of hidden layers, the number of nodes, the number of noise injections ($n$) and the size of the observation set ($N$). However, for high values of number of noise injections and large observations set, training networks with more than a single hidden layers and a large number of nodes is computationally demanding. For the $M$-$R$ data, the best accuracy we can reach is $0.889\pm 0.001$ with 2 hidden layers of 215 and 200 nodes respectively, for a dataset with $n=300$ and $N=30$; while for the $M$-$R$-$k_{2}$ the best accuracy we obtain is $0.878\pm 0.003$ for a single hidden layer network with 270 nodes, with the dataset corresponding to $n=300$ and $N=30$. We stress that, while a classification model is useful to distinguish between the two different low-density EOSs from which we built our dataset, this does not allow for a detailed analysis on the EOS properties in such regime, where uncertainties can still be large. We intend to improve this analysis in a future work by either allowing for a wider range of low-density EOSs from those commonly used in the literature, or by extending the analysis of the regression model to take into account this region as well.
@@
Let us now focus on the regression model, which aims to determine the remaining 15 parameters of the EOS. As in the previous case, we tested different neural networks varying the same parameters as above. We developed our models using the Python library Keras with TensorFlow. For every network we tested, we only allowed for a single hidden layer, while varying its number of nodes. We have seen that including more layers does not improve the efficiency of the model, while it requires more computational time. We also chose the ReLU activation function, with the He normal distribution as kernel initializer. The output layer consists of 15 nodes, corresponding to the 15 EOS parameters (7 $\rho$, 7 $c_{s}$, and $\Lambda$), together with the sigmoid activation function. The latter ensures that the speed of sound is bounded in the causal range $c_{s}\in(0,1)$  [^3] Note that, as a result of the renormalization process discussed in Sec. 3.1, all parameters, not only the speed of sound, are bounded in the interval $(0,1)$.. We use the Adam fitting algorithm [73], and we employed the msle loss function, that is
$l_{\tt msle}(y,\hat{y})=\frac{1}{N}\sum_{i=1}^{N}[\text{log}(1+y_{i})-\text{log}(1+\hat{y}_{i})]^{2},$ (3.4)
@@
Figure 2: A specific random sample for the speed-of-sound vs normalized mass-density parametrization, where $\rho_{f}=12\,\rho_{0}$ is the last fixed value for the mass density. The black line describes the real model, while blue curves represent all predicted profiles by the regression deep network applied on test dataset augmented by injecting Gaussian noise $n=100$ times, following the procedure described in Sec. 3.1. We also show the $2\,\sigma$ confidence interval for the network predictions as a blue shaded region, obtained by computing the mean and the standard deviation of the predictions at each value of renormalized mass density. The network predictions prefers to ‘settle’ around a mean value of $c_{s}$ such that $\langle c_{s}\rangle\sim 0.5$ for $\rho/\rho_{f}\to 1$. This behavior is retrieved generically, and is not related to the specific speed-of-sound model we show here.
@@
Figure 3: Predicted mean speed-of-sound $\langle c_{s}\rangle$ with a $3\,\sigma$ confidence interval as a function of the “real” (predicted from simulations) mean value for $c_{s}$ for all possible different speed-of-sound realizations in the test dataset. Almost all predicted values are not in agreement with the expected ones. The deep network tends to smooth out the sound speed profile, and misses large increases or decreases in the profile, which are responsible for oscillations on the mean value of $c_{s}$. Indeed, one can notice that the former are bound in the interval $(0.5,0.6)$, whereas for the latter one finds $\langle c_{s}\rangle_{\text{real}}\in(0.4,0.8)$. One can notice that deviations increase at the boundaries of such interval, where the mean value of the real speed-of-sound is either lower or higher to the predicted ones due to the presence of large oscillations in the real profile not reproduced by the network’s realization.
Let us now focus on the results for the predicted vacuum energy shift. In Fig. 4, we show the predicted mean $\Lambda$ with a $2\,\sigma$ confidence interval as a function of the real one for all $10$ possible values considered. We considered here the $n=100$ noise injections dataset. While the network can successfully predict a very large negative value of $\Lambda$, it tends to perform poorly as $\Lambda$ increases. While most of predicted values lie within the $2\,\sigma$ intervals of the “real” (predicted by simulations) ones, the measures are also consistent with $\Lambda=0$. In particular, we notice that the highest possible value for the vacuum energy shift, i.e. $\Lambda=(194\,\text{MeV})^{4}$, is not correctly predicted, even when considering a $3\,\sigma$ interval. We expect that the reason why our networks can better identify a large negative phase transition, rather than a (large) positive one, lies on the different properties of these two cases. EOSs with a vacuum energy phase transition can be divided into two classes [47]: EOSs that admits a wide range of values for $\Lambda$ and EOSs that are allowed only if a large negative jump in vacuum energy is introduced. These lead to different $M$-$R$ and $k_{2}$-$R$ relation properties. In particular, EOSs that admit a large negative $\Lambda$ display peculiar $M$-$R$ relations and the network can better distinguish them between all possible configurations. Moreover, our training dataset is biased towards lower values of the vacuum energy shift, since these values yield a larger set of physically acceptable solutions of the gravity equations. As a result, our network cannot confidently constrain the value of the jump in vacuum energy for the first family of EOSs.
@@
Figure 5: Two specific EOS models (black lines) and their predictions for the test dataset augmented by injecting Gaussian noise $n=10$ times, following the procedure described in Sec. 3.1. On the left (right) panel we show an EOS built from the AP4 (SLy) and with a negative (positive) jump in vacuum energy, with $\Lambda=-(194\,\text{MeV})^{4}$ ($\Lambda=(120\,\text{MeV})^{4}$). The two cases also use different speed-of-sound parametrizations. In almost all cases, the realizations have been predicted with the correct low-density model. At higher density scales, for the case of a negative vacuum energy transition, all predictions correctly display a sudden decrease in energy density. For a positive jump in $\Lambda$, however, the typical phase transition behavior is lost, and the energy density only slowly increases. The reason why our network can better identify a large negative vacuum energy, rather than a (large) positive one, lies on the different properties of these two cases. As it was pointed out in Ref. [47], there are two families of EOSs: those that admits a wide range of values for $\Lambda$ and those that are allowed only if a large negative jump in vacuum energy is introduced. This leads to different stellar properties. In particular, the second family displays peculiar $M$-$R$ relations and the network can better distinguish them between all possible configurations. Note also that our training dataset is biased towards lower values of the vacuum energy jump, since these choices lead to a larger set of physically acceptable solutions of the gravity equations.
@@
We now focus on the results coming from both the classification and the regression networks. This allows us to construct the full EOS, composed by the three regions described in section 2. In Fig. 5, we show two specific EOS models (black line) and their predictions for the $n=10$ test dataset. Note that the two expected EOSs correspond to different low density descriptions, different speed-of-sound parametrization and different vacuum energy shifts. On the left panel we present an EOS built from the AP4 and with a negative jump in vacuum energy, with $\Lambda=-(194\,\text{MeV})^{4}$. For the low density region, $8$ out of $10$ realizations have been predicted with the correct model (AP4). Moreover, at higher density scales all predictions correctly follow the decrease in energy density due to the vacuum energy shifts. On the right panel we show the case of an EOS built from the SLy model with a positive vacuum energy jump, with $\Lambda=(120\,\text{MeV})^{4}$. In this instance, only a single realization is predicted with the wrong low-density description. However, at high densities, the predictions follow the expected behavior less rigorously, in particular the increase in energy density caused by the positive jump in vacuum energy is less abrupt, and one loses the typical phase transition behavior. Once again, the reason why our networks can better identify a large negative phase transition, rather than a (large) positive one, lies on our training dataset being biased towards lower values of the vacuum energy shift and on the fact that these two cases belong to different EOS’ families, which leads to different stellar properties [47].
@@
This can be seen in Fig. 6, where we show the $M$-$R$ relations obtained by solving the TOV equations for both the expected EOS (black line) and the predicted ones for the two EOS models considered in Fig. 5. Note that the color mapping is maintained, i.e. an EOS described by a specific color line in Fig. 5 yields a $M$-$R$ curve of the same color in Fig. 6. We also include, as an example, a single set of observation points $O=\{M_{i},R_{i}\}$ with injected Gaussian noise as described in Eq. (3.2) with a $1\,\sigma$ confidence interval. For both cases presented, all of the observation points, which we remind are used as input for our networks, are within $3\,\sigma$ from the $M$-$R$ curves obtained from the predicted realizations. This is true even when the model predicts a wrong low-density model.
@@
Figure 6: Mass-radius relations for two specific EOS models, corresponding to those presented in Fig. 5. On the left (right) panel we show an EOS built from the AP4 (SLy) and with a negative (positive) jump in vacuum energy, with $\Lambda=-(194\,\text{MeV})^{4}$ ($\Lambda=(120\,\text{MeV})^{4}$). Curves obtained from solving the TOV equations for the real EOS are shown in black, whereas the network predictions are shown with different colors, maintaining the same mapping as in Fig. 5. We also include, as an example, a single set of observation points with injected Gaussian noise as described with a $1\,\sigma$ confidence interval, as described in Sec. 3.1, used as an input for the network. All of the observation points are within $3\,\sigma$ from the $M$-$R$ curves obtained from the predicted realizations.
@@
Finally, we discuss the $k_{2}$-$M$ relations obtained by solving Eq. 2.7 for the expected EOS (black line) and those obtained from our network. Once again, we maintain the color mapping as in Fig. 5. We do not include a set of observation points, to not complicate further the figure, however all mock observations are in agreement with the results produced from the predicted realizations. In Fig. 8, we show the fractional error for the tidal Love number $(k_{2}-k_{2,\,\text{real}})/k_{2,\,\text{real}}$. We can see that, typically, this is around $10\%$ at intermediate masses, while it can increase up to $30\%$ in the case of more massive stars.
@@
Figure 8: Fractional error for the tidal Love number as a function of the total stellar mass for two specific EOS models, corresponding to those presented in Fig. 5. On the right (left) panel we show an EOS built from the AP4 (SLy) and with a negative (positive) jump in vacuum energy, with $\Lambda=-(194\,\text{MeV})^{4}$ ($\Lambda=(120\,\text{MeV})^{4}$). We maintain the color mapping as in Fig. 5. The fractional error is typically around $10\%$ at intermediate masses, while it can increase up to $30\%$ in the case of more massive stars. This is to be expected since more massive configurations have higher central pressure, where the EOSs predictions deviates more significantly from the expected one, as can be seen from Fig. 5.
3.3 Bayesian deep network
Figure 9: A statistical sample of the predictions (features) as outputted from our Bayesian (probabilistic) deep network, and for two input choices according to our dataset’s numbering, corresponding to the $100$-th (right) and $1000$-th (left) input data point respectively. For a given input of masses, radii and tidal Love numbers, or equivalently a given input of the EOS, the output of the deep network corresponds to a particular realisation of the $7$ pairs of mass density-sound speed. We use the dataset where each input data point comprises of $30$ triplets of total mass, radius, tidal number, and we consider an $n_{\text{samples}}=100$ realisations of the output (coloured points). The red crosses denote the expected (“real”) values on the mass density-sound speed plane for the given EOS. The black dots and the associated error-bars are computed as the average and $2\sigma$ errors respectively, given the statistical sample for each column of the output’s array. As long as the predictions (red crosses) lie within the respective $2\sigma$ error bar, the prediction is considered successful, according to our evaluation pipeline, explained in Sec. 3.3. Notice that in the plots we chose not to discard predictions which are outside the physical range for the mass and sound speed we consider, i.e points lying outside the range $[0,1]$. Though individual predictions may lie outside this range, central values and respective error bars are consistent with our assumed physical range of values. It is also to be noticed how the size of error bars change with the assumed EOS provided as input to the deep network.
@@ 3.3 Bayesian deep network
$p({\boldsymbol{\theta}}|D)=\frac{p(\boldsymbol{\theta})\cdot p(D|\boldsymbol{\theta})}{\sum_{\boldsymbol{\theta}}p(D|\boldsymbol{\theta})\cdot p(\boldsymbol{\theta})}\equiv\frac{p(\boldsymbol{\theta},D)}{p(D)},$ (3.5)
@@
$q({\boldsymbol{\theta}})\approx p({\boldsymbol{\theta}}|D).$ (3.6)
@@
$\text{KL}\left(q||p\right)\equiv\sum_{\boldsymbol{\theta}}q(\boldsymbol{\theta})\log\frac{q(\boldsymbol{\theta})}{p(\boldsymbol{\theta}|D)}.$ (3.7)
In our approach, the approximate posterior, $q(\boldsymbol{\theta})$, will be based on some analytical, a priori ansatz, which we will compute after minimisation of the KL divergence (3.7). The resulting posterior function will then be a proxy to the true one in Bayes’ rule  [^4] We notice here that the KL divergence extends to the case of a continuous random variable through the replacement of the sum with an integral.. It can be shown that minimising the KL divergence amounts to maximising the quantity
$\mathds{E}_{q}\left[\log p(\boldsymbol{\theta},D)-\log q(\boldsymbol{\theta})\right]=\log p(\boldsymbol{\theta})-\text{KL}\left(q||p\right),$ (3.8)
@@
We implement our Bayesian deep network on TensorFlow making use of the class of DenseVariational layers which is part of the tensorflow_probability library. These are the probabilistic extension of the more standard Dense layers. We choose a LeakyReLU for our activation function and the mean-square error for our loss function respectively. The activation function is applied to all hidden layers, but not in the output layer - we comment further below about this choice. The minimisation of the loss function is performed through the RMSprop algorithm, which shows good-enough convergence at training. We found that a network with $4$ hidden layers, each with $16$ units, was a sufficiently minimal choice yielding satisfactory results for our purposes. Notice that the input and output of the network remained the same as with the standard (non-probabilistic) version of the deep network.
@@
$\bar{\rho}^{(J)}=\frac{1}{N}\sum_{i=1}^{i=N}\rho_{i}^{(J)},\;\;\;\;\bar{c}_{s}^{(J)}=\frac{1}{N}\sum_{i=1}^{i=N}c_{s\,i}^{(J)},\;\;\;J=1,\ldots,7,$ (3.9)
@@
$\bar{c_{s}}^{(J)}-2\sigma_{c_{s}}^{(J)}\leq c_{s\,\text{true}}^{(J)}\leq\bar{c_{s}}^{(J)}+2\sigma_{c_{s}}^{(J)}$ (3.10)
An example result of this process for a given data input is shown in Fig. 9 and 10 for a specific choice of input. For each data point we define as the accuracy $(\text{acc}_{i})\equiv N_{\text{accepted}}/7$, and the total accuracy as $\text{acc}\equiv\sum_{i}^{N}(\text{acc})_{i}/N,$ with $N$ the number of data points. The choice of the $2\sigma$ range for accepting/rejecting points is clearly a choice that we make as a reasonable credibility interval.
@@
$\text{mean accuracy}\pm\sigma\simeq\left(79\pm 16\right)\%.$ (3.11)
The spread of $16\%$ around the mean accuracy is somewhat large, implying that our model performs better on some EOS than others. It is to be emphasised that the features of our training/testing dataset, i.e the $7$ values of the sound speed, were drawn randomly from a normal distribution, as explained in Sec. 3.1. Our probabilistic network, therefore, did not have sufficient information to reconstruct the posterior functions more accurately and in turn, achieve a better accuracy. We also note that the accuracy of the prediction decreases with decreasing the number of input observations in the network, i.e the number of stellar mass, stellar radius and tidal Love number triplets. This we have verified trying with a dataset where each input to the network consisted of $15$ and $30$ pairs respectively. It is to be further noted that, the accuracy of the network’s predictions on test data may increase with decreasing training accuracy. This is because the error bars of output statistical samples will be larger at lower training accuracy, a situation which is clearly undesirable. We found that an optimal choice for the training accuracy is around $70\%$, which is a fair compromise between the previous issue, and the also undesirable problem of overfitting. At this point, we should emphasize that the quoted training accuracy is according to the way that TensorFlow translates the value of the loss function to an accuracy value. In our Bayesian setting, however, the accuracy cannot be derived by simply comparing the network’s single prediction with the true values of the features, but through a statistical analysis of the respective output samples.